### v1

# Delta-Doped Diamond via in-situ Plasma-Distance Control

Abstract
We present an approach for the CVD growth of diamond, where the sample is placed in a defined distance from the reactor baseplate, to which the plasma couples. We observe two previously unknown growth regimes. In the first case, the sample is positioned within three to five millimeters of the plasma, leading to a decreased growth rate, compared to a position inside the plasma and, additionally, to an increased nitrogen incorporation, allowing the fabrication of delta-doped layers with a thickness below 30 nm. In another regime, where the sample is more than 10 mm away from the plasma, no growth is observed. Instead, we assume a deposition of nitrogen-rich species on the diamond surface, which is incorporated during the growth of the following layer. All fabricated layers show NV emission, where the intensity correlates with the nitrogen incorporation. The growth techniques could allow the fabrication of highly doped thin films for quantum sensing applications, as well as layers with low NV concentration, for quantum computing. The new approaches are applicable not only for nitrogen incorporation but also for other defects, for example, phosphorus, which could open up new avenues for diamond-based electronics.
preprint: AIP/123-QED
The nitrogen vacancy (NV) center in diamond is a promising candidate for the implementation of quantum technologies, such as quantum sensors and quantum computers Prawer and Greentree (2008); Abobeih et al. (2022); Degen, Reinhard, and Cappellaro (2017); Taylor et al. (2008). The strong photoluminescence (PL) of dense NV layers, as well as the high coherence time of the electron spin, allow for highly sensitive sensing devices Herbschleb et al. (2019). The NV-center also offers applications in the field of quantum computing, where proximate nuclear spins can be used as additional qubits Bradley et al. (2019).
These applications place strong requirements on the diamond host material, including a high structural quality and low concentration of unintentional defects Bauch et al. (2020). One method for fabricating such diamond layers is via chemical vapor deposition (CVD). This technique allows the fabrication of high-quality diamond with a variety of dopants Schwander and Partes (2011); Kalish (1999).
One way to grow high quality diamond via CVD-growth, is through an enclosed sample holder, where the surface of the sample is on a lower level than the surface of the holder Ren et al. (2024); Mokuno et al. (2005); Nad, Gu, and Asmussen (2015). The depth of this enclosed area is usually a few mm Mokuno et al. (2005). With this technique, the structural quality of thick diamond layers, fabricated at high growth rates, has been improved with respect to samples, that were grown on an open holder  Mokuno et al. (2005).
However, some applications like quantum sensors require thin, so-called delta-doped diamond layers Ohno et al. (2012); Jaffe et al. (2020). The sensitivity of such a sensor can be enhanced by high signal and close proximity to the specimen through vertical positioning of the centers Jaffe et al. (2020). Moreover, the coherence time could also be improved in layers of lower dimensionality, compared to bulk material Schätzle et al. (2024a).
The fabrication of such thin films is challenging, especially since the growth rate of diamond increases drastically when nitrogen is introduced during CVD growth Jin and Moustakas (1994). Moreover, the incorporation efficiency of dopants like nitrogen and phosphorus in diamond is low Lobaev et al. (2017); Schätzle et al. (2023); Samlenski et al. (1995); Tallaire et al. (2006a). Various techniques have been proposed to grow nitrogen-doped diamond at low rates, allowing for a highly controlled thickness of the layers Ohno et al. (2012); Butler et al. (2017); Vikharev et al. (2016); Bogdanov et al. (2020, 2021); Kato et al. (2007). These include growth at low microwave powers Ohno et al. (2012) and subsequent electron irradiation Ohno et al. (2012), as well as surface termination and subsequent overgrowth Jaffe et al. (2020). Additionally, changes to the diamond reactor itself have been proposed, e.g., a laminar flow reactor, where the gas switching is more abrupt Bogdanov et al. (2020). Based on that, delta-doped layers with a high nitrogen concentration of more than 1000 ppm or NV concentration down to the single-center level have been obtained in high-quality diamond Jaffe et al. (2020); Ohno et al. (2012).
In this work, we present an alternative approach for the fabrication of nitrogen delta-doped layers in diamond via chemical vapor deposition. The method involves keeping the diamond sample in an adjustable retention position during growth, which is beneficial to the quality of the fabricated diamond Mokuno et al. (2005). This method offers control over the growth rate while maintaining a constant plasma composition. Adjustment of layer thickness and nitrogen incorporation can be achieved by changing the retention position with respect to the plasma. Using this technique, we can fabricate delta-doped layers with a thickness of a few nanometers and a high nitrogen incorporation.
Figure 1: Illustration of the different positions of the sample (2) with respect to the baseplate (1). The sample size is not to scale. a) The sample holder lies on the level of the baseplate, with the sample in the plasma (reference position). b) The sample holder is far away from the baseplate level, interrupting the diamond growth process. c) Intermediate positions, where the growth is influenced by the exact position of the sample, d) Image of a sample in the reactor in the reference position, as shown in a).
To fabricate said layers, a chemical vapor deposition diamond reactor was used, which is described elsewhere Schätzle et al. (2023). The main part of this reactor is a substrate lift that allows the sample to be removed and reintroduced to the plasma ball. The term plasma ball refers to the visible portion of the plasma that couples to the molybdenum baseplate. This allows for thin film growth with sharp interfaces, as the growth process is interrupted if the distance to the plasma is sufficiently high. To prevent an excessive cooling of the sample, when removed from the plasma, a heater that is situated within the sample holder was used. The heater keeps the sample at an elevated temperature, although the exact temperature of the sample cannot be determined in the retention position, because there is no line of sight to the attached pyrometer and the thermocouple attached to the heater is too far from the sample. In this work, we investigate the growth behavior and dopant incorporation in the intermediate regime, where the distance to the plasma ball is small enough to allow plasma species to reach the sample and react, as shown in Figure 1. The position is defined as the distance $d$ from the upper level of the base plate to which the plasma couples. In the following, we call the position where the holder is on the same level as the baseplate the "reference position".
The diamond growth was performed using hydrogen and methane as the main source gases. The total gas flow was 800 sccm at a pressure of 130 mbar, and the C/H-ratio is 1.75 % during doped growth. 15N was used as a dopant, with an N/C-ratio of 71,000 ppm. Under these conditions, a stack of six doped layers on (001)-oriented CVD diamond substrates (Element6, electronic grade) was grown. The sample was placed on a flat holder with a diameter of 2.54 cm. Using 12C-enriched source gas, we can distinguish the grown, doped layers from the substrate below. Between the doped layers, we fabricated intrinsic layers which have a 13C concentration of 1.1 %, which is the natural abundance. Growth was performed on a first sample (S1), varying the distance to the baseplate, to which the plasma couples, between 20 and 0.1 mm during the 35-minute doped growth. The following intrinsic layer was grown for 10 min in the adapted position and subsequently 5 min in the reference position. The growth conditions for one of the doped doped layers and the following buffer layer are summarized in Table 1, and a more detailed description of the growth processes and sample preparation can be found in the Supplementary Material.
Table 1: Growth conditions during the fabrication of one of the doped layers and the following buffer layer of sample S1.
Step CH4-conc. 13C-conc. N/C position time
[%] [% of CH4] [ppm] [$m$] [min]
1 1.75 0.0 71000 20 / 15 / 10 / 35
5 / 3 / 0.1
2 2.0 1.1 0 20 / 15 / 10 / 10
5 / 3 / 0.1
3 2.0 1.1 0 0 5
To assess the doping of diamond, that was grown in the varied growth positions, time-of-flight secondary ion mass spectrometry (ToF-SIMS) measurements were performed. Fig. 2 shows the layer stack of sample S1.
The measurements revealed a strong change of the growth rate and nitrogen incorporation when changing the position of the sample during growth. We distinguish between three different growth modes. The first one is the trivial one, where the sample holder is only 0.1 mm below the baseplate level, where the growth conditions are established and well knownSchätzle et al. (2023). The second position is slightly below this first position at a distance of 3 mm, where a significant drop of the growth rate and an increase of the nitrogen incorporation was observed. The third regime starts at a distance of more than 15 mm. Here, the incorporation of nitrogen is still observed, but no carbon incorporation is indicated by the ToF-SIMS profile, as shown in Fig. 2. It should be noted, that the observed, distance-dependent behavior may be influenced by the reactor geometry and especially by the diameter of the opening in the baseplate.
In a position 0.1 mm below the baseplate level, a layer with a thickness of 172 nm was obtained, determined from the full width at half maximum of the nitrogen concentration. This corresponds to a growth rate of 295 nm/h. The layer is also indicated by the dip in the 13C-concentration. The temperature during growth was 730 °C, as determined by a pyrometer. The nitrogen concentration in the layer is on average 22 ppm, corresponding to an incorporation efficiency of $3.1\cdot 10^{-4}$, in line with literature results Lobaev et al. (2017); Schätzle et al. (2023); Samlenski et al. (1995); Tallaire et al. (2006a).
In lower positions, which are still close to the plasma (3 to 5 mm), a significant change in the growth behavior is observed. The growth rate is drastically reduced to 45 nm/h, with a layer thickness of 26.5 nm in a distance $d$ of 3 nm. We also observe a high nitrogen incorporation, with a peak concentration of 228 ppm, resulting in an incorporation efficiency of $3.2\cdot 10^{-3}$. These effects could be related to a lower substrate temperature in these lower positions, which generally favors nitrogen incorporation Lobaev et al. (2017); Yan and Vohra (1999); Tallaire et al. (2006b). Additionally, the lower position might be protected from direct exposure to the plasma and thus allow less reactive species to reach the sample, lowering the growth rate. This change in the growth regime is beneficial for the fabrication of delta-doped layers because the low growth rate allows the adjustment of the layer thickness with high precision. Furthermore, the high nitrogen concentration could allow a high performance of fabricated sensors, due to the high NV concentrations Teraji et al. (2024).
The first layer in Fig. 2, which was fabricated at a distance of 20 mm from the top level of the baseplate, shows that a significant amount of nitrogen was incorporated (blue curve), even though the sample was not in contact with the plasma. This suggests that the reactive plasma species are mobile and have a sufficient lifetime outside the visible plasma ball to reach the sample 2 cm below. We also observe a high incorporation of nitrogen in these growth conditions, reaching a peak concentration of 15 ppm. More interestingly, no dip in the 13C concentration of this layer is observed, which suggests that no significant diamond growth happens at this depth.
We assume that the nitrogen in this layer was incorporated during the subsequent intrinsic growth process. While nitrogen is in the plasma, it could be deposited on the diamond surface and terminate it. We assume that the conditions do not allow for diamond growth and that the nitrogen is incorporated during the growth of the subsequent layer. In literature, it has been shown that the overgrowth of a nitrogen-terminated surface can lead to the incorporation of nitrogen in the growing diamond Jaffe et al. (2020).
Figure 2: ToF-SIMS profile of sample S1, showing the layer stack of six doped layers (blue). The distance $d$ to the base plate was varied in each of the grown layers (20, 10, 0.1, 15, 5, 3 mm), starting from the substrate. The layers are additionally marked by the dip in the 13C concentration (black).
To check this assumption, we grew the same six-layer stack again, but kept the position at 20 mm distance and varied the process time from 1 min to 35 min. The resulting depth profile is shown in Fig. 3. There is no significant or linear increase of the layer thickness, indicated by the increase in nitrogen concentration, with respect to the growth time. Additionally, the nitrogen concentration increases with increasing growth time, hinting that there is no constant growth and defect incorporation, which would be independent of time, but rather a deposition of nitrogen species. Furthermore, again, no significant dip in the 13C concentration is observed.
This growth concept is in opposition to the general assumption that the sample must be in contact with or very close to the plasma to allow single crystal diamond growth Gicquel et al. (2001); Mankelevich and May (2008); Mokuno et al. (2005); Muchnikov et al. (2010). In fact, we observe diamond growth, indicated by the change in isotopic composition, up to 10 mm from the top level of the baseplate.
The different growth behavior opens up new avenues for diamond epitaxy. They allow for changes in the growth rate and defect incorporation in doped layers, while keeping constant plasma conditions. Furthermore, the deposition of defect species in positions that are more than 10 mm away from the base plate could enable new techniques for color center fabrication.
We also used this alternative growth mode to incorporate other species, like phosphorus. There, we again observed a high concentration spike, one order of magnitude higher than the incorporation in the reference position, where the sample is in the plasma. The associated data is shown in the Supplementary Material.
Figure 3: ToF-SIMS profile of sample S2, where the growth time in a position was varied (35, 10, 5, 1, 2, 20 min from substrate). No dip in the 13C concentration is observed (black). The nitrogen incorporation in the layers (blue) increases linearly with increasing growth time, as shown in the inset.
To assess the suitability of the fabricated layers for quantum-technological applications, the PL intensity of single layers that were grown in the different positions was measured. We chose the same growth conditions as before, but reduced the N/C-fraction in the gas phase to 10,400 ppm, to reduce the nitrogen concentration in the thin films. The positions with respect to the baseplate were 0.1 mm (S3), 3 mm (S4), and 20 mm (S5). From previous processes in the reference position, we estimate the P1 center concentration in sample S3 to 2 ppm.
We then performed PL measurements, using a Renishaw InVia confocal microscope, as described in the Supplementary Material. We acquired depth slices within a 20 µm x 20 µm section of the sample, as shown in Fig. 4 a) to c). The stepsize in x- and z-direction was 1 µm and 0.5 µm, respectively. To compensate for fluctuations of the laser power or optical losses, the signal acquired at the zero phonon line of the NV- center (637 nm) was normalized with respect to the Raman peak at 573 nm.
Figure 4: Example spectra from bright spots of the PL-slices, shown in a) to c). a) Depth-slice through sample S3, grown 0.1 mm below the reference position. b) and c) Depth-slice through sample S4 and S5, where the doped layer was grown 3 mm and 20 mm below te baseplate level. The color scale on the right side is valid for all plots, a) to c).
The sample that was grown 0.1 mm below the reference position shows strong PL intensity with a clearly visible zero phonon line of the NV center. Additionally, SiV- emission, with a zero phonon line at 737 nm Tóth et al. (2015) was observed. Silicon is often incorporated during CVD diamond growth and could be introduced to the reactor via plasma-etching of the glass bell jar Yang et al. (2024). We assume that the high Si-incorporation arises from surface damage caused by etching processes during plasma changes, especially when the sample is close to the reference position. These etching processes could lead to the formation of facets with enhanced defect incorporation Ashfold et al. (2020).
The sample, grown in a position 3 mm below, shows a similar PL intensity, although the layer thickness is expected to be nearly 7 times lower. This most likely arises from the high nitrogen incorporation and confirms that this is a suitable way to obtain thin films with a high NV concentration. An example spectrum, as shown in Fig. 4, does not reveal the presence of a high concentration of SiV centers. However, the curve suggests the presence of both NV- and NV0, with a zero phonon line at 575 nm Manson et al. (2013).
The layer that was grown 20 mm below the reference position has a thickness of only 14 nm, as well as the lowest nitrogen incorporation in the six-layer stack. As expected, the PL intensity of sample S5 is low compared to the other samples. The single graph in Fig. 4 reveals a slight bump barely above the noise level, indicating the presence of only a few NV centers.
The coherence times of sample S4 were determined, using a home-built setup for measuring NV ensembles, as described in the Supplementary Material. We obtain T2* values of 0.38 µs and T2 is 7.1 µs. We assume, the values are mainly limited by the high nitrogen concentration. The measured T2-time implies a P1-center concentration of 23 ppm Bauch et al. (2020). The layer grown at roughly 7 times the nitrogen concentration in the gas phase exhibits a nitrogen concentration of 228 ppm, as determined by ToF-SIMS. This would suggest a nitrogen concentration of 33 ppm in sample S4, if the incorporation efficiency was the same. The relatively higher coherence time could arise from a changing incorporation efficiency or the quasi two-dimensional nature of the layer Schätzle et al. (2024b); Bogdanov et al. (2021).
In this work, we present a new diamond growth technique that allows the fabrication of nitrogen delta-doped layers with high nitrogen concentration. We show two distinct growth regimes that strongly differ from established processes. The photoluminescence measurements hint at a high NV-concentration in the layers, fabricated in a retention position, 3 mm below the reactor baseplate, making them suitable for use in quantum sensing applications. Moreover, we observe a smaller nitrogen incorporation during growth, 20 mm below the plasma, leading to the formation of small ensembles of NVs. These layers with a low NV concentration could be beneficial in the field of quantum computing.
The new approaches may open up new avenues for diamond growth in various other areas. For example, enhanced phosphorus incorporation could be advantageous for the fabrication of diamond-based electronic devices. We are currently studying the surface of the diamond after a process where the sample is 20 mm below the reference position to gain a better understanding of the behavior and possible applications.
This work was supported by the european union, via the projects SPINUS (Grant No. 101135699) and AMADEUS (Grant No. 101080136).